Use set builder notation to give a description of each of these sets.

a. {0, 2, 4, 6, 8,10,12,14,16}

b. {−3,−2,−1, 0, 1, 2, 3}
1
Expert's answer
2021-09-30T02:16:54-0400

Let us use the set builder notation to give a description of the following sets.


a. "\{0, 2, 4, 6, 8,10,12,14,16\}=\{n\in\Z:0\le n\le 16\text{ and }n\text{ is even}\}."


b. "\{−3,−2,−1, 0, 1, 2, 3\}=\{n\in\Z: |n|\le 3\}."